OpenStudy (anonymous):

The area of the trapezoid is 24 and the bases are 3 and 5. Find the height. please help

OpenStudy (anonymous):

\[A=\frac{b_{1}+b_{2}}{2}*h\]\[24=\frac{5+3}{2}*h\]\[24=\frac{8}{2}*h\]\[24=4*h\]\[\frac{24}{4}=h\]\[6=h\]